• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kuchina

    Log entry time 08:00:35 on October 9,2006

    Entry number 186656

    keyword=Owl Shift Summary

    Start shift with run 1158 in setting A10 started during the previous shift.

    Run 1158 5.0M LH2 Sieve in HRS-L Raster ON
    Run 1159 5.0M LH2 Sieve in HRS-L Raster ON
    Run 1160 5.0M LH2 Sieve in HRS-L Raster ON
    Run 1161 2.3M LH2 Sieve in HRS-L Raster ON

    2:03am Collimators switched to left arm open, right arm sieve

    Run 1162 - probably junk

    2:20am Asked for 50 uA

    Run 1163 303K LH2 Sieve in HRS-R Raster ON

    3:05am Changed target to Optics, asked for 80 uA and raster off

    Run 1164 21K Optics Sieve in HRS-R Raster OFF

    Slow rates, changed prescale factors to ps1=1, ps3=14

    Run 1165 2.3M Optics Sieve in HRS-R Raster OFF

    4:25am Changed to Carbon single foil target

    Run 1166 0.9M Carbon Sieve in HRS-R Raster OFF

    5:26am Beam is off till 5:33

    Run 1167 Junk: IOC16 lost communication,

    To restore HRS-R dipole we followed the procedure:
    Reboot IOC16,
    Reboot NMR,
    Typed in p0,
    on HRS Right Dipole NMR GUI selected:
    Mode=MAN, Field Pol=NEG, Search=ON
    Communication back.

    6:15am Asked for Controlled access for the cryotarget people.

    7:00am David Meekins is here. Cryotarget work started.
    7:30am Target warm up is complete.